OpenStudy (pinklion23):

@jigglypuff314 @welshfella

OpenStudy (pinklion23):

Jose received $30 for his birthday and put it in his piggy bank. Each week, he puts 4 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks he has saved 4 dollars. His sister Marisol makes a table of reasonable values for w. Marisol selects 16 as a value for w. What is the corresponding table entry for D? A. 34 B. 434 C. 50 D. 94

OpenStudy (welshfella):

Ok D = 4w + 30 is the formula for working out the amount of money in the bank after w weeks So you need to replace w in this formula by 16 and calculate D. Can you try this?

OpenStudy (pinklion23):

sure thing

OpenStudy (pinklion23):

94 :) that was my guess tbh

OpenStudy (pinklion23):

am i right?

OpenStudy (pinklion23):

?

OpenStudy (welshfella):

4*16 + 30 = 64 + 30 = 94 Yes

OpenStudy (pinklion23):

thx i have more if you dont mind i will post new question so you guys get more medals :)